OpenStudy (anonymous):

HELP PLEASE I WILL GIVE METALS AND BECOME A FAN choose an equation for the line that is parallel to the given line that passes through the given point for equations 1 and 2. 1. y = 5x + 10; (2,14) a. y= 1/5x - 4 b. y= -1/5x -4 c. y= 5x - 68 d. Y= 5x +4

OpenStudy (anonymous):

2. Y = 1/5X - 10 ; (15,-16) a. y= -5x +19 b. y= 1/5x + 91/5 c. y= 1/5x -19 d. y= 5x -19

OpenStudy (anonymous):

If we want have parallel lines their slope(m) should be equal . (if you want understand it draw two line with a slope ) So what is the slope for each lines ?;) (Find the m(slope) in first line ;) ) @yolo131

OpenStudy (anonymous):

tell whether the lines for each pair of equations are parallel, perpendicular, or neither for questions 3 through 4. I do care I was just starting to type the next question I will draw the lines in just a sec

OpenStudy (anonymous):



OpenStudy (anonymous):

you just want answer ?!!!!!!

OpenStudy (anonymous):

sorry I suck at math

OpenStudy (anonymous):

no I want to understand it to

OpenStudy (anonymous):

its just hard for me

OpenStudy (anonymous):

I meant was nt that ...! if you want to understand a thing should try for that . i said you ... tell me : What is the slope in the first equation ?!

OpenStudy (anonymous):

if you cant answer this cant answer this question to ...

OpenStudy (anonymous):

3. y = -7/8x -1 and -8x + 7y = -21 a. parallel b. perpendicular c. neither can you answer this one?

OpenStudy (anonymous):

please step by step ... wait to get answer in the first then other caz they are same questions ... that is nt helpful for you ...

OpenStudy (anonymous):

4. y = -1/4x + 10 and -2x + 8y = 6 a. parallel b. perpendicular c. neither

OpenStudy (anonymous):

....

OpenStudy (anonymous):

can you post the answers? i would like to learn but it is better for me to see it then it be explained

OpenStudy (anonymous):

im sorry i suck at math :( let me see if my friend can help you

OpenStudy (anonymous):

@EMoneh97 can you help ?

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Ok first since these answers are completely random. @yolo131 , have you figured out any of the answers yet?

OpenStudy (anonymous):

no...... I hate and suck at all math

OpenStudy (anonymous):

i think yolo 131 must be pay atention !;) ok ? u ready !?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

ill do my best

OpenStudy (anonymous):

ok ! thank you ! so first the first question . ok ?;)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

choose an equation for the line that is parallel to the given line that passes through the given point for equations 1 and 2. 1. y = 5x + 10; (2,14) a. y= 1/5x - 4 b. y= -1/5x -4 c. y= 5x - 68 d. Y= 5x +4

OpenStudy (anonymous):

So ... to start . can u tell me what is the slope (m) in the first question ?;) ( the number before x .!;) )

OpenStudy (anonymous):

1. 1 y =5x+10 Compare this equation with y=mx+b slope m = 5 The slope of a line parallel to the above line will be the same m= 5,point (2, 14) Find b by plugging the values of m & the point in y=mx+b 14= 10+b b=4 m=5 Plug value of the slope and b in y = mx +b The required equation is y =5x+4

OpenStudy (anonymous):

y = 5x + 10 What is the slope there ?

OpenStudy (anonymous):

@EMoneh97 :Thanks !;)

OpenStudy (anonymous):

5